Solve the folling 2 x 2 systems using substitutions
y= 3x + 12
y= -x +4

Answer by elima(1433)

y= 3x + 12
y= -x +4
plug -x + 4 into the first equation for the y;
-x + 4 = 3x +12
-x -3x +4 = 12
-4x = 12 - 4
-4x = 8
x = -2
now plug that in the equation for x;
y = 3(-2) + 12
y = -6+12
y = 6
:)